# Open-Source ESP32-C3 Wi-Fi Drivers for Static Analysis

### v2

@@ 4.4. Implementation
The ESP32-Open-MAC project lists the registers that are necessary to achieve basic transmission and reception on the ESP32. In order to identify the equivalent registers on the ESP32-C3, we build a sketch that performs simple UDP transmission and analyze it using Ghidra’s (Agency, 2019) decompiler. We also add volatile memory sections for the peripheral ranges identified in previous reverse engineering efforts (lcgamboa, 2023). The author also provides a fork of Espressif’s QEMU that adds support for Wi-Fi functionality on the ESP32-C3. We modify this fork to log all peripheral accesses.
We replace only the external function names in the binaries with function names of similar length. Furthermore, we declare these functions in the main ESP-IDF sketch to the original function names in the object files. The resulting sketch is both loaded to the emulator and flashed to an ESP32-C3-Devkit-M1.
@@
Some important registers and their deduced purpose are shown in Table 1.
@@
Setting up interrupts on the ESP32-C3 is well documented. However, the Wi-Fi interrupt is not exposed to developers. Through decompilation, we find that the Wi-Fi driver uses CPU interrupt number 1. To register a custom Wi-Fi interrupt handler, we find that first the source register corresponding to interrupt 1 (0x600c2000) has to be cleared. The interrupt has to be disabled by masking bit 1 of 0x600c2104. The custom interrupt handler can be set using intr_handler_set provided by ESP-IDF. Finally, the interrupt is enabled again by setting bit 1 of 0x600c2104.
@@
We implement the driver using the architecture described in Figure 2. We use an open AP, as the encryption functionality of Wi-Fi peripheral hardware is not fully reverse-engineered. For initialization and connecting to an AP, we use the API calls esp_wifi_init and esp_wifi_start provided by ESP-IDF.
These calls initialize and configure the Wi-Fi peripheral, set the internal interrupt handler for Wi-Fi, and initiate a connection to the AP. The successful completion of these tasks can be monitored through a system notification WIFI_EVENT_STA_CONNECTED. Once the notification is received, we kill the Wi-Fi task using the undocumented pp_post function and set up the custom interrupt handler. Finally, our driver and MAC tasks are started. We use PfIP (Vogelgesang et al., 2025) as the IP stack due to its intermittency-aware execution.
Static Analysis
For static analysis, we use the Platin framework (Maroun et al., 2024). Platin uses a modified version of the clang compiler that collects program flow information. The compiler generates a metadata file from the program flow information and transforms it into an ILP, which, in turn, can be solved using programs such as lp_solve (Berkelaar et al., 2004) or gurobi (Gurobi Optimization, LLC, 2023). For the purpose of this paper, we annotate the current consumption for each state in the device’s energy model based on the datasheet (Espressif Systems, 2022). Additionally, Platin distinguishes between the energy penalty of switching on and off any given energy state (wcec-turn-on and wcec-turn-off) and the energy consumption of a device in the state energy-stay-on. wcec-always-on represents the energy required when the device is in any given state. In our model, we set both the switching on and off penalties to zero and model the device-switching behavior pessimistically.
@@ 5. Evaluation
This section presents the evaluation of our open-source Wi-Fi driver. Section 5.1 describes the hardware and network setup as well as the task system used for the experiments. Additionally, we detail the required assumptions for the static resource-consumption analysis using the Platin analysis framework (Maroun et al., 2024). In Section 5.2, we then examine and discuss the results from the static analysis.
@@ 5.1. Evaluation Setup
To validate the functionality of the Wi-Fi driver, we perform an experiment with three tasks. The mac_task interacts with the PfIP network stack, the hw_task initiates the transmission of packets required by the mac_task and handles receiving, and a test_task sends out a packet using the PfIP stack and waits for an answer from other devices in the network. We validated the functionality of our open-source driver with Wireshark captures of the network traffic.
Static Analysis
To analyze the worst-case resource demand of our Wi-Fi driver, we use the Platin analysis framework (Maroun et al., 2024) in combination with the device-model analysis of the WoCA approach (Raffeck et al., 2024). The focus of our evaluation lies solely on the driver functions, independent of the runtime and operating system. While a whole-system analysis, which includes the operating system, is generally possible (Schuster et al., 2019), we exclude the FreeRTOS components and the interaction with the FreeRTOS scheduler from our analysis. Likewise, we exclude the callbacks into the PfIP network stack from the analysis.
We perform the worst-case resource-demand analysis both for individual driver functions and for a task comprising multiple device interactions. For each individual driver function, we determine an upper bound on the execution time. The evaluation task transmits a packet and actively waits for the completion of the transmission. After starting the transmission, the Wi-Fi transceiver transmits the packet and waits for an acknowledgment (ACK). Once the transmission is completed, the transmission returns to standby mode. We derive an upper bound on both the time and energy demand with respect to the different device states of the Wi-Fi transceiver during the task execution.
@@ 5.2. Evaluation Results
Most functions of our open-source Wi-Fi driver perform only operations on memory-mapped registers to control the transceiver and exhibit relatively low execution time and energy consumption. wifi_mac_handle_rx copies the payload data of the received package between buffers, so the analysis has to assume the worst case of the maximum packet size. Similarly, the analysis has to assume the worst-case number of packets to handle for the execution time of wifi_handle_rx and the maximum time until an ACK is received for the execution time of wifi_wait_for_tx. Finally, the difference between the always-on and device-aware WCEC bounds underlines the importance of considering the states of the Wi-Fi transceiver in a fine-granular manner.
@@ 6. Related Work
Typical approaches for battery-free wireless communication use Bluetooth Low Energy (BLE), LoRa , or backscatter-based wireless communications (Cai et al., 2023; Talla et al., 2017; Jiang et al., 2023). Intermittency-aware operation has been successfully demonstrated, especially for BLE and LoRa.
@@
Wireless networking is a key challenge in Battery-free IoT designs due to intermittency. Therefore, the networking stack deployed on these devices must be capable of intermittent execution. (Fu et al., 2023) proposed a fully intermittency-aware networking stack. Schoeberl et al. developed tpIP, a time-predictable IP stack for cyber-physical networks (Schoeberl and Pedersen, 2018). tpIP is designed to allow static analysis through bounded loops and avoids dynamic memory allocation. They also estimated WCET for the stack’s operations. Similarly, Vogelgesang et al. designed PfIP, a transactional networking stack with worst-case bounds for every transaction (Vogelgesang et al., 2025). These approaches are agnostic to the underlying physical layer.
De Winkel et al. introduced a battery-free smartwatch running FreeBie (de Winkel et al., 2022), a bidirectional BLE stack for intermittently-powered Bluetooth devices. The device stored checkpoints for all processes when power failure was imminent and restored them when power was available. Additionally, due to intermittency, two devices may not be awake at the same time for communication. In (Geissdoerfer and Zimmerling, 2021, 2022), these challenges are addressed through probabilistic modeling of latency and charge times, respectively.
Similarly, a LoRa-based battery-free image sensing system that adaptively configures its transmission mode based on the available energy was developed in (Nardello et al., 2019). Raffeck et al. demonstrated a LoRa-based device using WoCA, a whole system worst-case analysis approach for intermittent wireless devices (Raffeck et al., 2024). For static analysis, Platin was used with loop bounds calculated based on the LoRa device configuration.
Greentooth is a proposed low-power wireless protocol designed for battery-free devices (Babatunde et al., 2024). Greentooth synchronizes communication between two devices using time division multiple access (TDMA)-style scheduling and low-power wake-up radios.
Wi-Fi remains under-explored due to its higher power requirement compared to BLE or LoRa. The ubiquity of Wi-Fi networking applications and availability of off-the-shelf inexpensive hardware such as the ESP32-C3 promises battery-free innovation in the field of, for example, smart-home applications, where an existing Wi-Fi network is usually available.
@@
Open-source drivers exist for several Wi-Fi network interface cards from, for example, Atheros or Mediatek (Project, 2024). These require no binary blobs and are often supported by the manufacturers themselves. However, they require a Linux kernel and PCI-E hardware support and are generally not intended for low-power SoCs.
The OpenWiFi project (Jiao et al., 2020) developed an open-source FPGA-based software-defined radio for the 802.11 standard paired with an RF frontend. However, FPGAs are not commonly used in consumer hardware and the platform is intended to be a tool for researchers.
On low-power devices, the ZeroWi project (Bentham, 2020) focuses on reverse-engineering the capabilities of the Cypress CYW43438 Wi-Fi SoC. The CYW43438 consists of an internal ARM Cortex CPU, which is programmed using non-free firmware provided by Cypress. Drivers written for this chip then send commands in the SDIO format over SPI to communicate with the firmware.
For the Espressif products, early attempts at reverse engineering were performed by Uri Shaked, who used it as part of his Wokwi simulator for ESP32 (Shaked, 2019). The QEMU-based open-source PICSimLab simulator introduces support for Wi-Fi emulation on the ESP32 and ESP32-C3 devices (lcgamboa, 2022). Based on these attempts, the ESP32-Open-MAC project (Devreker, 2023) aims to fully open-source the Wi-Fi driver on the ESP32. Our work not only ported ESP32-Open-MAC onto the RISC-V–based ESP32-C3 but also integrated it with (1) a resource-consumption model for static analysis and (2) a transactional network stack that is suitable for battery-free devices.
@@ 7. Future Work & Conclusion
The current energy model does not consider packet reception. Therefore, future work needs to consider a more detailed device model: That is, pessimism in the bounds can be reduced by a more fine-grained device model. For example, in , the peripheral is not actively transmitting after the end of transmission until ACK is received. Additionally, energy penalties for switching on/off energy states could be accurately considered. Further, we use a static peripheral configuration. The resource-consumption bounds are impacted by peripheral configurations, such as data rate of transmission. An interesting avenue for future work is to identify optimal configurations under energy constraints. The ESP32-Open-MAC project also has recently introduced initialization independent of FreeRTOS. Therefore, an OS-aware implementation could also result in easier static analysis, for example, by exploiting Rust’s statically analyzable async/await pattern.
In this work, we presented a fully open-source Wi-Fi driver for the ESP32-C3, based on the ESP32-Open-MAC project. We successfully port the project from the Tensilica-based ESP32 to the RISC-V–based ESP32-C3 using common reverse-engineering techniques. Finally, we present a proof-of-concept for performing static resource-consumption analysis on the physical operations of the Wi-Fi peripheral using the extended Platin worst-case analysis toolkit. For the future, we envision more exchange between the static-analysis and the security communities. More interdisciplinary research on this topic could avoid a future of insecure and unsafe battery-free IoT systems.
[ backgroundcolor=green!10, linecolor=green!70,linewidth=2pt,topline=false,bottomline=false, innertopmargin=0pt, innerbottommargin=5pt, innerleftmargin=6pt, innerrightmargin=6pt, frametitleaboveskip=5pt, frametitlefont=, frametitlealignment=, frametitle=The implementation is available at: https://git.nt.uni-saarland.de/open-access/pfip ]
@@ Acknowledgements.
This work is supported by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) as part of SPP 2378 (Resilient Worlds: project number 502615015, Resilient Power-Constrained Embedded Communication Terminals, ResPECT). We would also like to thank all ESP32-Open-MAC contributors for their valuable contributions to the community.References
@@
- Abella et al. (2015) Jaume Abella, Carles Hernández, Eduardo Quiñones, Francisco J. Cazorla, Philippa Ryan Conmy, Mikel Azkarate-askasua, Jon Pérez, Enrico Mezzetti, and Tullio Vardanega. 2015. WCET analysis methods: Pitfalls and challenges on their trustworthiness. In Proceedings of the 10th International Symposium on Industrial Embedded Systems (SIES ’15). IEEE, 39–48. https://doi.org/10.1109/SIES.2015.7185039
@@
- Babatunde et al. (2024) Simeon Babatunde, Arwa Alsubhi, Josiah D. Hester, and Jacob Sorber. 2024. Greentooth: Robust and Energy Efficient Wireless Networking for Batteryless Devices. ACM Transactions on Sensor Networks 20, 3 (may 2024), 66:1–66:31. https://doi.org/10.1145/3649221
@@
- Cai et al. (2023) Zhipeng Cai, Quan Chen, Tuo Shi, Tongxin Zhu, Kunyi Chen, and Yingshu Li. 2023. Battery-Free Wireless Sensor Networks: A Comprehensive Survey. IEEE Internet of Things Journal 10, 6, March 15 (mar 2023), 5543–5570. https://doi.org/10.1109/JIOT.2022.3222386
- Choi et al. (2022) Jongouk Choi, Larry Kittinger, Qingrui Liu, and Changhee Jung. 2022. Compiler-Directed High-Performance Intermittent Computation with Power Failure Immunity. In Proceedings of the 28th Real-Time and Embedded Technology and Applications Symposium (RTAS ’22). IEEE, 40–54. https://doi.org/10.1109/RTAS54340.2022.00012
- Cypress Semiconductor Corporation (2021) Cypress Semiconductor Corporation 2021. CYW43439 Single-Chip IEEE 802.11 b/g/n MAC/Baseband/Radio with Integrated Bluetooth 5.2 Compliance. Cypress Semiconductor Corporation. https://www.mouser.com/datasheet/2/196/Infineon_CYW43439_DataSheet_v03_00_EN-3074791.pdf
- de Winkel et al. (2022) Jasper de Winkel, Haozhe Tang, and Przemyslaw Pawelczak. 2022. Intermittently-Powered Bluetooth That Works. In Proceedings of the 20th Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’22) (Portland, Oregon) (MobiSys ’22). ACM, New York, NY, USA, 287–301. https://doi.org/10.1145/3498361.3538934
- Dengler et al. (2023) Eva Dengler, Phillip Raffeck, Simon Schuster, and Peter Wägemann. 2023. FusionClock: Energy-Optimal Clock-Tree Reconfigurations for Energy-Constrained Real-Time Systems. In Proceedings of the 35th Euromicro Conference on Real-Time Systems (ECRTS ’23), Vol. 262. 6:1–6:24. https://drops.dagstuhl.de/opus/volltexte/2023/18035/pdf/LIPIcs-ECRTS-2023-6.pdf
@@
- Espressif Systems (2022) Espressif Systems 2022. ESP32-C3 Series Datasheet. Espressif Systems. https://www.espressif.com/sites/default/files/documentation/esp32-c3_datasheet_en.pdf
@@
- Geissdoerfer and Zimmerling (2021) Kai Geissdoerfer and Marco Zimmerling. 2021. Bootstrapping Battery-free Wireless Networks: Efficient Neighbor Discovery and Synchronization in the Face of Intermittency. In 18th USENIX Symposium on Networked Systems Design and Implementation (NSDI 21). USENIX Association, 439–455. https://www.usenix.org/conference/nsdi21/presentation/geissdoerfer
- Geissdoerfer and Zimmerling (2022) Kai Geissdoerfer and Marco Zimmerling. 2022. Learning to Communicate Effectively Between Battery-free Devices. In 19th USENIX Symposium on Networked Systems Design and Implementation (NSDI 22). USENIX Association, 419–435. https://www.usenix.org/conference/nsdi22/presentation/geissdoerfer
- Gurobi Optimization, LLC (2023) Gurobi Optimization, LLC. 2023. Reference Manual. gurobi.com/.
@@
- Jiao et al. (2020) Xianjun Jiao, Wei Liu, Michael T. Mehari, Muhammad Aslam, and Ingrid Moerman. 2020. openwifi: a free and open-source IEEE802.11 SDR implementation on SoC. In 91st IEEE Vehicular Technology Conference, VTC Spring 2020, Antwerp, Belgium, May 25-28, 2020. IEEE, IEEE, 1–2. https://doi.org/10.1109/VTC2020-Spring48590.2020.9128614
@@
- Maroun et al. (2024) Emad Jacob Maroun, Eva Dengler, Christian Dietrich, Stefan Hepp, Henriette Herzog, Benedikt Huber, Jens Knoop, Daniel Wiltsche-Prokesch, Peter P. Puschner, Phillip Raffeck, Martin Schoeberl, Simon Schuster, and Peter Wägemann. 2024. The Platin Multi-Target Worst-Case Analysis Tool. In Proceedings of the 22nd International Workshop on Worst-Case Execution Time Analysis (WCET ’24), Vol. 121. Schloss Dagstuhl - Leibniz-Zentrum für Informatik, 2:1–2:14. https://doi.org/10.4230/OASIcs.WCET.2024.2
- Mottola et al. (2024) Luca Mottola, Arslan Hameed, and Thiemo Voigt. 2024. Energy Attacks in the Battery-less Internet of Things: Directions for the Future. In Proceedings of the 17th European Workshop on Systems Security (EuroSec ’24). ACM, 29–36. https://doi.org/10.1145/3642974.3652283
- Nardello et al. (2019) Matteo Nardello, Harsh Desai, Davide Brunelli, and Brandon Lucia. 2019. Camaroptera: A batteryless long-range remote visual sensing system. In Proceedings of the 7th International Workshop on Energy Harvesting & Energy-Neutral Sensing Systems (EnsSys ’19). ACM, 8–14. https://doi.org/10.1145/3362053.3363491
@@
- Raffeck et al. (2024) Phillip Raffeck, Johannes Maier, and Peter Wägemann. 2024. WoCA: Avoiding Intermittent Execution in Embedded Systems by Worst-Case Analyses with Device States. In Proceedings of the 25th ACM SIGPLAN/SIGBED International Conference on Languages, Compilers, and Tools for Embedded Systems (LCTES ’24). ACM, 83–94. https://doi.org/10.1145/3652032.3657569

### v3

@@ 4.4. Implementation
The ESP32-Open-MAC project lists the necessary MMIO registers to achieve basic transmission and reception on the ESP32. On the ESP32, the MMIO addresses belong to the 0x3ff73000-0x3ff75fff range, but on the ESP32-C3 are mapped to 0x60033000-0x60035fff. The offsets of ESP32-C3’s MMIO registers from the base address differ from those of the ESP32. In order to identify the equivalent registers on the ESP32-C3, we write a firmware that performs simple UDP transmission and analyze it using Ghidra’s (Agency, 2019) decompiler. In Ghidra, we mark the memory ranges of the peripheral identified with QEMU (lcgamboa, 2023) as volatile and identify the functions that access these memory locations. Functions with prefixes such as hal_mac_* or wdev_mac_* were found to handle MAC/PHY layer operations, which are of particular interest for our reverse engineering.
Among the identified functions, we hook into those that are listed as undefined symbols in object files within the libpp.a archive of the proprietary blobs. The resulting firmware with the hooks is loaded into our QEMU fork.
@@
An excerpt of the most important registers identified through our reverse engineering is shown in Table 1, along with the deduced purpose of these registers.
@@
The details of the Wi-Fi interrupt handling are not exposed to developers. Through decompilation, we find that the Wi-Fi driver uses CPU interrupt number 1, which is marked as reserved in the manual (Espressif Systems, 2025b), and uses an undocumented register for that interrupt source at address 0x600c2000. With this information, registering a Wi-Fi interrupt handler follows the usual procedure on the ESP32-C3. First, the (undocumented) interrupt source register (0x600c2000) is cleared, and the interrupt is disabled by clearing bit 1 of the (documented) interrupt-enable register (0x600c2104). The interrupt handler is registered using ESP-IDF’s intr_handler_set. Finally, the interrupt is enabled by setting bit 1 of 0x600c2104 and loading the interrupt number to 0x600c2000.
@@
We implement the driver using the architecture shown in Figure 1. We use an open AP, as the encryption functionality of the Wi-Fi hardware is not fully reverse-engineered. For initialization and connecting to an AP, we use ESP-IDF’s API. The functions esp_wifi_init and esp_wifi_start initialize and configure the Wi-Fi, set the internal interrupt handler for Wi-Fi, and initiate an AP connection. Their successful completion can be monitored through the system notification WIFI_EVENT_STA_CONNECTED. Once the notification is received, the Wi-Fi task can be killed using the undocumented pp_post function, which queues instruction codes to the Wi-Fi task started by the driver. Finally, our driver and MAC tasks are started. We use PfIP (Vogelgesang et al., 2025) as the IP stack due to its intermittency-aware execution.
@@ 5. Evaluation
We perform a static resource analysis of our Wi-Fi driver to assess the ability to bound the program flow and resource demand. Section 5.1 describes the hardware and network setup as well as the task system used for the experiments. Additionally, we detail the required assumptions for the static resource-consumption analysis using the Platin analysis framework (Maroun et al., 2024). In Section 5.2, we examine and discuss the analysis results.
@@ 5.1. Evaluation Setup
To validate the driver’s functionality, we write a firmware that sends UDP packets through PfIP’s network stack. PfIP is provided with the mac_task’s transmit handle, which in turn queues the packet for physical transmission by the hw_task. The responses from the network are passed back through PfIP’s receive callback. We capture the network traffic using Wireshark to monitor our driver’s functional correctness.
Static Resource Analysis
To analyze the worst-case resource demand of our Wi-Fi driver, we use the Platin analysis framework (Maroun et al., 2024). With the device-model analysis of the WoCA approach (Raffeck et al., 2024), Platin supports device-aware resource-consumption analysis for the ESP32-C3. Platin uses a modified version of the clang compiler that collects user-annotated knowledge, derives metainformation about the program flow, and tracks the relation between source and object level regarding this control-flow information. Using this metainformation and hardware-specific cost models, Platin formulates the costs of and constraints on the program behavior as an ILP suitable for common mathematical solvers, such as lp_solve (Berkelaar et al., 2004) or Gurobi (Gurobi Optimization, LLC, 2025), to obtain the maximum resource demand. For the purpose of this paper, we annotate the current consumption for each state in the device’s energy model based on the values provided by the manufacturer in the datasheet (Espressif Systems, 2022). For the processor, we rely on the model for the ESP32-C3 already present in Platin.
The focus of our evaluation lies on the driver functions, independent of the runtime and operating system. While a whole-system analysis, which includes the operating system, is generally possible (Schuster et al., 2019), we exclude the FreeRTOS components and callbacks into the PfIP network stack from our analysis for this paper.
We analyze benchmarks including individual driver functions and an evaluation task comprising multiple device interactions. For each driver function, we determine an upper bound on the execution time. The evaluation task sends a packet and actively waits for the transmission’s completion. After starting the transmission, the Wi-Fi transceiver transmits the packet and waits for an acknowledgment (ACK). Once the transmission is completed, the transmission returns to standby mode. We derive an upper bound on both the time and energy demand with respect to the different device states of the Wi-Fi transceiver during the task execution.
@@ 5.2. Evaluation Results
Most functions of our open-source Wi-Fi driver perform only operations on memory-mapped registers to control the transceiver and exhibit a relatively low execution time and energy consumption. The few notably higher resource bounds relate to worst-case analysis assumptions for the control flow. wifi_mac_handle_rx moves the payload of the received package between buffers, so the analysis has to assume the worst case of the maximum packet size. Similarly, the analysis has to assume the worst-case number of packets for the execution time of wifi_handle_rx and the maximum time until an ACK is received for wifi_wait_for_tx.
The difference between the always-on and device-aware WCEC bound underlines the importance of considering the states of the Wi-Fi transceiver in a fine-granular manner. We demonstrated that the reverse engineering of our driver code enables deriving the necessary control-flow bounds and performing accurate static resource analysis of Wi-Fi operations on the ESP32-C3.
@@ 6. Related Work
Wireless networking is a key challenge in batteryless designs due to intermittency. Networking stacks deployed on such devices must be fully aware of intermittent execution (Fu et al., 2023). In contrast to our work on Wi-Fi, other approaches for batteryless wireless communication use Bluetooth Low Energy (BLE), LoRa, or backscatter-based wireless communications (Cai et al., 2023; Talla et al., 2017; Jiang et al., 2023). As a time-predictable IP stack, tpIP (Schoeberl and Pedersen, 2018) is designed to allow static analysis through bounded loops and avoids dynamic memory allocation. Similarly, Vogelgesang et al. designed PfIP, a networking stack with transactional semantics and worst-case bounds for every transaction (Vogelgesang et al., 2025). Transactional semantics break down the overall task into atomic sub-tasks (or transactions) that can be independent of each other. These approaches are agnostic to the underlying physical layer.
FreeBie (de Winkel et al., 2022) is a bidirectional BLE stack designed for intermittent communication, demonstrated with a batteryless smartwatch. Cameroptera (Nardello et al., 2019) is a LoRa-based batteryless image sensing system that adaptively configures its transmission mode based on the available energy. RockClimb (Choi et al., 2022) uses WCET bounds without device awareness to guarantee the reaching of checkpoints. Yarahmadi et al. (Yarahmadi and Rohou, 2020) employ WCEC-aware checkpoint placement without considering the resource demand of peripheral devices. WoCA introduces device-aware whole-system worst-case analysis for intermittent systems, including a model for LoRa-based communication (Raffeck et al., 2024).
Without an intermittency-aware protocol, two devices may not be awake at the same time for communication. Greentooth (Babatunde et al., 2024) synchronizes communication between two devices using time division multiple access (TDMA)-style scheduling and low-power wake-up radios. Probabilistic modeling of latency or charge times (Geissdoerfer and Zimmerling, 2021, 2022) likewise achieves synchronized communication.
For intermittent devices, Wi-Fi is under-explored due to its high power demand but promises batteryless innovation, for example, in fields like smart-home applications.
@@
Outside the world of low-power SoCs, open-source drivers exist for several Wi-Fi network interface cards, often supported by the manufacturers themselves, for example, from Atheros or Mediatek (Project, 2024). The OpenWiFi project (Jiao et al., 2020) developed an open-source FPGA-based software-defined radio for the 802.11 standard paired with an RF frontend.
On low-power devices, the ZeroWi project (Bentham, 2020) reverse-engineers the capabilities of the Cypress CYW43438 Wi-Fi SoC, which is programmed using non-free firmware. Drivers written for this chip then send commands in the SDIO format over SPI to communicate with the firmware. Early attempts at reverse engineering Espressif’s Wi-Fi driver were performed by Uri Shaked as part of the Wokwi simulator for ESP32 (Shaked, 2019). The QEMU-based open-source PICSimLab simulator introduces support for Wi-Fi emulation on the ESP32 and ESP32-C3 (lcgamboa, 2022). Based on these attempts, the ESP32-Open-MAC project (Devreker, 2023) aims to fully open-source the Wi-Fi driver on the ESP32. Our work not only ported ESP32-Open-MAC onto the RISC-V–based ESP32-C3 but also integrated it with a resource-consumption model for static analysis and a transactional network stack that is suitable for batteryless devices.
@@ 7. Future Work & Conclusion
The use of Wi-Fi for intermittently-powered batteryless systems is an ongoing research challenge since network protocols are not designed with intermittent operation in mind (Ahmed et al., 2024). With this work, we contribute to this challenge with our reverse-engineered and publicly-available Wi-Fi drivers, along with their static worst-case energy-consumption analysis. Our analysis results allow intermittent systems to execute with runtime guarantees on the power-failure–free execution of operations.
Future work needs to consider a more detailed device model to reduce the pessimism in the bounds. For example, in , the peripheral is not actively transmitting after the end of transmission until the ACK is received. Additionally, energy penalties for switching on/off energy states could be accurately considered. We also use a static peripheral configuration. The resource-consumption bounds are impacted by peripheral configurations, such as data rate of transmission. Future work could identify optimal configurations under energy constraints. ESP32-Open-MAC has recently introduced an initialization of the Wi-Fi peripheral independent of FreeRTOS. Therefore, an OS-aware implementation could also result in easier static analysis, for example, by exploiting Rust’s statically analyzable async/await pattern.
In this work, we presented a fully open-source Wi-Fi driver for the ESP32-C3, based on the ESP32-Open-MAC project. We successfully port the project from the Xtensa-based ESP32 to the RISC-V–based ESP32-C3 using common reverse-engineering techniques. Finally, we present a proof-of-concept for performing static resource-consumption analysis on the Wi-Fi peripheral’s physical operations using the extended Platin worst-case analysis toolkit.
[ backgroundcolor=tolerableGreen!25, linecolor=tolerableGreen,linewidth=2pt,topline=false,bottomline=false, innertopmargin=0pt, innerbottommargin=3pt, innerleftmargin=6pt, innerrightmargin=6pt, frametitleaboveskip=3pt, frametitlefont=, frametitlealignment=, frametitle=The implementation is available at: ] https://git.nt.uni-saarland.de/open-access/pfip
@@ Acknowledgements.
This work is supported by the German Research Foundation (DFG) as part of SPP 2378 (Resilient Worlds: project number 502615015, Resilient Power-Constrained Embedded Communication Terminals, ResPECT) and the NGI0 Core Fund under grant agreement No 101092990. We would also like to thank all ESP32-Open-MAC contributors for their valuable contributions to the community.References
@@
- Abella et al. (2015) Jaume Abella, Carles Hernández, Eduardo Quiñones, Francisco J. Cazorla, Philippa Ryan Conmy, Mikel Azkarate-askasua, Jon Pérez, Enrico Mezzetti, and Tullio Vardanega. 2015. WCET analysis methods: Pitfalls and challenges on their trustworthiness. In Proceedings of the 10th International Symposium on Industrial Embedded Systems (SIES ’15). 39–48. https://doi.org/10.1109/SIES.2015.7185039
@@
- Babatunde et al. (2024) Simeon Babatunde, Arwa Alsubhi, Josiah D. Hester, and Jacob Sorber. 2024. Greentooth: Robust and Energy Efficient Wireless Networking for Batteryless Devices. ACM Transactions on Sensor Networks 20, 3 (2024), 66:1–66:31. https://doi.org/10.1145/3649221
@@
- Cai et al. (2023) Zhipeng Cai, Quan Chen, Tuo Shi, Tongxin Zhu, Kunyi Chen, and Yingshu Li. 2023. Battery-Free Wireless Sensor Networks: A Comprehensive Survey. IEEE Internet of Things Journal 10 (2023), 5543–5570. https://doi.org/10.1109/JIOT.2022.3222386
- Choi et al. (2022) Jongouk Choi, Larry Kittinger, Qingrui Liu, and Changhee Jung. 2022. Compiler-Directed High-Performance Intermittent Computation with Power Failure Immunity. In Proceedings of the 28th Real-Time and Embedded Technology and Applications Symposium (RTAS ’22). 40–54. https://doi.org/10.1109/RTAS54340.2022.00012
- de Winkel et al. (2022) Jasper de Winkel, Haozhe Tang, and Przemyslaw Pawelczak. 2022. Intermittently-Powered Bluetooth That Works. In Proceedings of the 20th Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’22) (Portland, Oregon) (MobiSys ’22). 287–301. https://doi.org/10.1145/3498361.3538934
- Dengler et al. (2023) Eva Dengler, Phillip Raffeck, Simon Schuster, and Peter Wägemann. 2023. FusionClock: Energy-Optimal Clock-Tree Reconfigurations for Energy-Constrained Real-Time Systems. In Proceedings of the 35th Euromicro Conference on Real-Time Systems (ECRTS ’23), Vol. 262. 6:1–6:24. https://doi.org/10.4230/DARTS.9.1.2
@@
- Espressif Systems (2022) Espressif Systems. 2022. ESP32-C3 Series Datasheet. https://www.espressif.com/sites/default/files/documentation/esp32-c3_datasheet_en.pdf
- Espressif Systems (2025a) Espressif Systems. 2025a. ESP IoT Development Framework. https://www.espressif.com/en/products/sdks/esp-idf
- Espressif Systems (2025b) Espressif Systems. 2025b. ESP32-C3 Technical Reference Manual. https://www.espressif.com/sites/default/files/documentation/esp32-c3_technical_reference_manual_en.pdf
@@
- Geissdoerfer and Zimmerling (2021) Kai Geissdoerfer and Marco Zimmerling. 2021. Bootstrapping Battery-free Wireless Networks: Efficient Neighbor Discovery and Synchronization in the Face of Intermittency. In 18th USENIX Symposium on Networked Systems Design and Implementation (NSDI 21). 439–455. https://www.usenix.org/conference/nsdi21/presentation/geissdoerfer
- Geissdoerfer and Zimmerling (2022) Kai Geissdoerfer and Marco Zimmerling. 2022. Learning to Communicate Effectively Between Battery-free Devices. In 19th USENIX Symposium on Networked Systems Design and Implementation (NSDI 22). 419–435. https://www.usenix.org/conference/nsdi22/presentation/geissdoerfer
- Gurobi Optimization, LLC (2025) Gurobi Optimization, LLC. 2025. Reference Manual. gurobi.com
@@
- Jiao et al. (2020) Xianjun Jiao, Wei Liu, Michael T. Mehari, Muhammad Aslam, and Ingrid Moerman. 2020. openwifi: a free and open-source IEEE802.11 SDR implementation on SoC. In Proceedings of the 91st IEEE Vehicular Technology Conference (VTC 2020). IEEE, 1–2. https://doi.org/10.1109/VTC2020-Spring48590.2020.9128614
@@
- Maroun et al. (2024) Emad Jacob Maroun, Eva Dengler, Christian Dietrich, Stefan Hepp, Henriette Herzog, Benedikt Huber, Jens Knoop, Daniel Wiltsche-Prokesch, Peter P. Puschner, Phillip Raffeck, Martin Schoeberl, Simon Schuster, and Peter Wägemann. 2024. The Platin Multi-Target Worst-Case Analysis Tool. In Proceedings of the 22nd International Workshop on Worst-Case Execution Time Analysis (WCET ’24), Vol. 121. 2:1–2:14. https://doi.org/10.4230/OASIcs.WCET.2024.2
- Mottola et al. (2024) Luca Mottola, Arslan Hameed, and Thiemo Voigt. 2024. Energy Attacks in the Battery-less Internet of Things: Directions for the Future. In Proceedings of the 17th European Workshop on Systems Security (EuroSec ’24). 29–36. https://doi.org/10.1145/3642974.3652283
- Nardello et al. (2019) Matteo Nardello, Harsh Desai, Davide Brunelli, and Brandon Lucia. 2019. Camaroptera: A batteryless long-range remote visual sensing system. In Proceedings of the 7th International Workshop on Energy Harvesting & Energy-Neutral Sensing Systems (EnsSys ’19). 8–14. https://doi.org/10.1145/3362053.3363491
@@
- Raffeck et al. (2024) Phillip Raffeck, Johannes Maier, and Peter Wägemann. 2024. WoCA: Avoiding Intermittent Execution in Embedded Systems by Worst-Case Analyses with Device States. In Proceedings of the 25th ACM SIGPLAN/SIGBED International Conference on Languages, Compilers, and Tools for Embedded Systems (LCTES ’24). 83–94. https://doi.org/10.1145/3652032.3657569